# Entanglement dynamics in intensity-dependent double Jaynes-Cummings model for squeezed coherent thermal states

### v6

@@ 3.1 Wigner function distributions for SCTS
The density operator $\hat{\rho}$ for SCTS is given in Eq. (1). Wigner functions $W(\alpha)$ for SCTS in IDDJCM are depicted in Fig. 4. In these plots $\bar{n}_{c}=2.0,\bar{n}_{s}=0.1$ and $\bar{n}_{th}=0.1$ are chosen. It is observed that at $t=0$, $W(\alpha)$ is positive and it is a shifted Gaussian; however, at $t=2$, the peak of $W(\alpha)$ decreases significantly and at some region it becomes negative. The negative peak is almost $30\%$ of the peak at $t=2$ After that, the negativity of $W(\alpha)$ decreases and at $t=6$ , it almost resembles the initial peak at $t=0$. As time increases further, it is noticed that negativity of Wigner function increases again at $t=8$ and again it decreases. So, it is observed that $W(\alpha)$ changing value from positive to negative and vice versa as time progresses. It indicates that the state of radiation oscillating between a classical state and non-classical state as an effect of atom-field interaction in IDDJCM.
@@
Figure 5 represents the Wigner function distributions for SCTS in DJCM. In this case also, $W(\alpha)$ starts from a shifted Gaussian shape at $t=0$. At $t=2$, it is positive peak comes down very drastically and in some region it becomes negative (almost $50\%$) of the highest of value. As time increases, the shape of $W(\alpha)$ changes and negative values decreases. However, it is contrasting compared to the case with IDDJCM that $W(\alpha)$ does not become positive again and mimic the initial state. The Wigner function for SCTS shows negative values for a longer period of time in DJCM. It shows that SCTS become more non-classical as a result of atom-field interaction in DJCM as compared to the case in IDDJCM.
4 Effects of single photon exchange interaction on the entanglement dynamics
Now we study the effects of single photon exchange between the two cavities on entanglement dynamics. Recently, there are studies, in which photons can hop from one atom-cavity system to another. Pandit et. al., [28, 31] have studied the effects of cavity-cavity interaction on the entanglement dynamics of a generalized double Jaynes-Cummings model. It is shown that for larger value of $\kappa$, entanglement for the atom-atom subsystem can be protected. They have also investigated the role of $\kappa$ in entanglement transfer between atom-atom and field-field subsystems. In another work, Laha [33] has investigated the effects of beam-splitter which is equivalent to photon exchange interaction on entanglement dynamics in DJCM for radiation states in vacuum, coherent and thermal states. The effects of photon exchange interaction is modeled by the following Hamiltonian:
@@ 4 Effects of single photon exchange interaction on the entanglement dynamics
where $\kappa$ is the cavity-cavity photon exchange coupling term. Without intensity dependence this Hamiltonian becomes
@@
Figure 6 shows the effects of cavity-cavity single photon exchange interaction on the entanglement dynamics for SCTS in IDDJCM. The strength of this cavity-cavity interaction via single photon exchange is characterized by the parameter $\kappa$. To study the effects of $\kappa$, $\bar{n}_{c}=2,\bar{n}_{s}=1$ and $\bar{n}_{th}=0.1$ are taken. From Fig. 6(a) it is observed that for $\kappa=0.1$, i.e., $\kappa<\lambda$, the peaks of atom A-atom B entanglement pulses decrease with time (described by blue curve); however, the lengths of ESDs do not change considerably. The field a-field b entanglement which is represented by the black curve shows different behaviour. If we compare Figs. 6(d) and 1(d), it can be seen that ESDs are removed from the dynamics and the amplitude of $N(t)$ increases with time. For both atom A-field a (green curve) and atom A-field b (red curve), the periodic nature of $N(t)$ gets destroyed; nevertheless, the amplitude of $N(t)$ does not change noticeably. For $\kappa=\lambda$ ($\kappa=1$), the peaks of $C(t)$ decrease rapidly with time and also the length of first ESD increases; however, the length of other ESDs decreases at some places. For field-field entanglement, the amplitude of $N(t)$ rises up sharply and then it becomes oscillatory. For the atom A-field a and atom A-field b, all ESDs get removed from the dynamics, though the amplitude of $N(t)$ in these subsystems are small. As we increase $\kappa$ further, i.e, $\kappa>\lambda$ ($\kappa=5$), it is observed that ESD in the dynamics of $C(t)$ are removed initially and many smaller peaks appear in the dynamics. From the plot, it is evident that this photon exchange interaction is creating atom-atom entanglement in the system. In the field-field entanglement, $N(t)$ shows reduction in its amplitude, which can be explained as the increase on number of smaller peaks and disappearance of ESDs in $C(t)$; the entanglement is transferred from $N(t)$ to $C(t)$. Also, the frequency of field-field entanglement increases significantly. The atom A-field a entanglement is decreased and it almost becomes zero at some places. $N(t)$ for atom A-field b does not change considerably; yet, it becomes zero, when atom-atom and field-field entanglements are maximum. As we increase the value of $\kappa$ to $10$, we have a situation where $\kappa>>\lambda$, i.e, photon exchange coupling between the cavities is more powerful than coupling between the atoms and radiation field. All the ESDs from the dynamics of $C(t)$ are removed and the value of $C(t)$ increases by a large amount. This increase in atom-atom entanglement makes the field-field entanglement to decrease considerably; the entanglement is transferred from field-field subsystem to atom-atom subsystem. $N(t)$ for atom A-field a and atom A-field b are almost zero everywhere except for some very small initial peaks. So, we observe that for large photon exchange coupling between the cavities, ESDs are removed from atom-atom dynamics and also the amplitude increases significantly. $N(t)$ for field-field entanglement too gets rid of ESDs with the addition of $\kappa$ in the system; however, its amplitude decreases with increasing value of $\kappa$.
@@
Figure 7 shows the effects of photon exchange interaction for SCTS in DJCM. The maximum value of of $N(t)$ for field-field subsystem increases by almost three times as a result of photon exchange interaction. For $\kappa=0.1$, which is a low value compared to the atom-field coupling constant $\lambda$. For this low value of $\kappa$, the lengths of ESDs for atom-atom entanglement increase and other smaller peaks disappear. The atom A-field a entanglement comes down significantly and at some places, $N(t)$ almost becomes zero. In the case of the atom A-field b entanglement, we observe that the ESDs are removed from the dynamics of $N(t)$. For $\kappa=1.0$, the length of ESD of $C(t)$ becomes larger, as small peaks disappear. $N(t)$ for field-field entanglement increases very sharply and then comes down again. Its maximum value also decreases as compared to the case when $\kappa=0.1$. In atom A-field a subsystem, ESD occurs as a result of increase in $\kappa$. No appreciable change is noticed in atom A-field a entanglement. As $\kappa$ is increased further to $5$, a drastic change is observed in the entanglement dynamics of $C(t)$. We see that all the ESDs are removed from the dynamics and its value also increases significantly at every instant of time. In field-field entanglement, we notice that the maximum value of $N(t)$ increases and the dynamics becomes very oscillatory. $N(t)$ for atom A-field a and atom A-field b shows ESD initially and after that it stays nonzero with very low amplitude. When $\kappa>>\lambda$, i.e., for $\kappa=10$, $C(t)$ increases more and remains high. We observe interesting patterns in the dynamics of $N(t)$ for field-field entanglement. In this case, wave packets of entanglement are formed and amplitude of wave packets increase significantly. $N(t)$ for atom A-field a and atom A-field b entanglements become zero. From this, we conclude that the photon exchange interaction creates entanglement in the atom-atom and field-field subsystems. However, in the case of IDDJCM, we observe that $C(t)$ increases and $N(t)$ for field-field entanglement decreases. On the other hand, in the case of DJCM, both $C(t)$ and $N(t)$ for the field-field entanglement increase. In both IDDJCM and DJCM, atom A-field a and atom A-field b entanglements become zero with increasing $\kappa$.
@@ 5 Effects of dipole-dipole interaction between the two atoms on entanglement
The effects of the dipole-dipole interaction on the entanglement dynamics for the IDDJCM is represented in Fig. 8. The effects for $g_{d}=0.1$ and $g_{d}=1.0$ are almost similar, as we observe in photon exchange interaction for $\kappa=0.1$ and $1.0$. For $g_{d}=5$, we observe that field-field entanglement is different from the dynamics for $\kappa=5$. In this case no oscillation is found in the dynamics, though its amplitude increases significantly. The behaviour of other entanglements are quite similar. When $g_{d}$ is increased to $10$, all the ESDs are removed from the dynamics of $C(t)$. Also, the amplitude of oscillations get increased; but, the oscillations die after sometime. For the field-field entanglement, it is noticed that the downward fall of the dynamics stops. The atom-field entanglements vanish for this high value of dipole-dipole interaction.
@@
The effects of dipole-dipole interaction for DJCM with SCTS is plotted in Fig. 9. From Fig. 9(a), it can be seen that all the dynamics remain almost unchanged for $g_{d}=0.1$. Only field-field entanglement rises up for large time, making ESDs disappear. For $g_{d}=1.0$, peaks of $C(t)$ come down remarkably; but, the length of ESDs remains same. The field-field entanglement increases more and all the ESDs are removed from the dynamics. The amplitude of $N(t)$ for atom A-field a comes down; however, $N(t)$ for the atom A-field b does not change considerably. For $g_{d}=5$, it is observed that new peaks begin to appear in the dynamics of $C(t)$ and all the ESDs get removed in the dynamics for $g_{d}=10$. The field-field entanglement remains high as before. However, the atom-field entanglements almost disappear for $g_{d}=10$.
@@ 6 Effects of spin-spin Ising interaction between the two atoms on entanglement
In the present section we study the effects of Ising-type interaction between the two atoms on the entanglement dynamics. The Hamiltonian with Ising-type interaction for IDDJCM and DJCM can be written as follows:
@@
The effects of Ising interaction on entanglement for IDDJCM are shown in Fig. 10. The plots show very different effects as are observed in DJCM for SCS and G-L states. From Fig. 10(a) and 10(b), it can be seen that for $J_{z}=0.1$ and $0.3$, we cannot distinguish the plots from the plots with $J_{z}=0$ for $C(t)$ and $N(t)$ (see Figs. 1 and 10). However, for atom A-field b entanglement, the amplitude of $N(t)$ comes down and ESDs are also present in the dynamics. In the case of field a-field b entanglement, for the above mentioned values of $J_{z}$, we observe that the ESDs are removed from the dynamics and amplitude increases. For $J_{z}=0.7$, small peaks appear along with the large peaks in the dynamics of $C(t)$ (red curve) and also the amplitude decreases. For atom A-field a subsystem, the amplitude of $N(t)$ comes down a little and ESDs are also removed in due course of time. In the case of atom A-field b, we do not see any noticeable change in the dynamics. However, for the field a-field b subsystem, the amplitude increases further, and no ESD is observed. For $J_{z}=1.0$, small peaks in the dynamics of $C(t)$ appear earlier in the dynamics and its height increases with time, while the initial peak decreases. As a result, the duration of ESDs gets smaller. For other subsystems, $N(t)$ follows the dynamics similar to that corresponding to $J_{z}=0.7$, viz., with decreased peak height for atom A-field a entanglement, while increased peak height for atom A-field b and field a-field b subsystems.
@@
The effects of Ising interaction on entanglements in DCJM are depicted in Fig. 11. In this case also, we observe that for small values of $J_{z}$ (0.1 and 0.3), $C(t)$ and $N(t)$ for all subsystems do not change considerably. Only field-field entanglement increases slightly for $J_{z}=0.3$. For $J_{z}=0.7$, length of ESD for $C(t)$ increases slightly, as the small bump disappears from the dynamics. The amplitudes of atom-field entanglement do not change noticeably; however, the length of ESD decreases for atom A-field b entanglement. For $J_{z}=1.0$, two small peaks appear in the dynamics of $C(t)$ which makes the length of ESDs smaller. All other entanglement dynamics remain almost unchanged. So, we see that Ising interaction is not much effective to remove sudden death of entanglement for atom-atom subsystem and also from the atom A-field b subsystem.

### v7

@@ 3.1 Wigner function distributions for SCTS
The density operator $\hat{\rho}$ for the squeezed coherent thermal state (SCTS) is given in Eq. (1). The corresponding Wigner function $W(\alpha)$, which provides a quasi-probability distribution in phase space, is used as a reliable indicator of the quantum state’s nonclassicality. In particular, the presence of negative regions in $W(\alpha)$ signals strong nonclassical behavior.
The Wigner function for IDDJCM and DJCM are presented in Figs. 3 and 4. The contrasting behavior between the two models highlights an important physical insight: the intensity-dependent coupling in IDDJCM plays a stabilizing role, enabling temporary nonclassicality, and also permitting the system to re-enter a classical regime periodically. In DJCM, the absence of intensity dependence leads to more persistent dephasing and deeper excursions into nonclassical territory. Hence, while both models generate nonclassical light due to atom–field interactions, IDDJCM exhibits quasi-periodic modulation of classicality, whereas DJCM drives the system more deeply and irreversibly into nonclassical states.
These findings suggest that tuning interaction parameters and the nature of the coupling allows control over the quantum-to-classical transition and the degree of nonclassicality in the radiation field, which is crucial for quantum technologies relying on engineered quantum states of light.
4 Effects of photon exchange interaction (PEI) on the entanglement dynamics
Now we study the effects of single photon exchange between the two cavities on entanglement dynamics. Recent studies have explored scenarios where photons can hop from one atom-cavity system to another. The effects of photon exchange interaction is modeled by the following Hamiltonian:
@@ 4 Effects of photon exchange interaction (PEI) on the entanglement dynamics
where $\kappa$ is the cavity-cavity photon exchange coupling term. In particular, Pandit et. al., [24, 27] have studied the effects of cavity-cavity interaction on the entanglement dynamics of a generalized double Jaynes-Cummings model. It is shown that for larger value of $\kappa$, entanglement for the atom-atom subsystem can be protected. They have also investigated the role of $\kappa$ in entanglement transfer between atom-atom and field-field subsystems. In another work, Laha [29] has investigated the effects of beam-splitter which is equivalent to photon exchange interaction on entanglement dynamics in DJCM for radiation states in vacuum, coherent and thermal states. Without intensity dependence this Hamiltonian becomes
@@
Figure 5 shows the effects of cavity-cavity single photon exchange interaction on the entanglement dynamics for SCTS in IDDJCM. The strength of this cavity-cavity interaction via single photon exchange is characterized by the parameter $\kappa$. To study the effects of $\kappa$, $\bar{n}_{c}=2,\bar{n}_{s}=1$ and $\bar{n}_{th}=0.1$ are taken. From Fig. 6(a), it is observed that for $\kappa=0.1$, i.e., $\kappa<\lambda$, the peaks of atom A-atom B entanglement pulses decrease with time (described by blue curve); however, the lengths of ESDs do not change considerably. The field a-field b entanglement which is represented by the black curve shows different behaviour. If we compare Figs. 6(d) and 1(d), it can be seen that ESDs are removed from the dynamics and the amplitude of $N(t)$ increases with time. For both atom A-field a (green curve) and atom A-field b (red curve), the periodic nature of $N(t)$ gets destroyed; nevertheless, the amplitude of $N(t)$ does not change noticeably. The PEI allows photons to hop between cavities, effectively coupling the two field modes and disrupting their independent evolution. This interaction mixes the Fock components across the cavities, increasing photon number uncertainty and breaking the regular structure essential for coherent rephasing. As a result, the system loses its well-defined Rabi frequencies tied to fixed photon numbers, leading to spectral irregularity and enhanced dephasing. The entanglement dynamics become more complex and less periodic, with entanglement flowing not only within each atom–field pair but also across atom–atom and field–field subsystems, resulting in a more irregular and intertwined evolution.
For $\kappa=\lambda$ (i.e., $\kappa=1$), the peaks of the atom-atom entanglement $C(t)$ decay rapidly over time, and the duration of the first episode of entanglement sudden death (ESD) increases. However, the durations of subsequent ESD intervals tend to decrease in certain regions. In the case of field-field entanglement, the measure $N(t)$ exhibits a sharp initial rise followed by oscillatory behavior. For the hybrid subsystems—atom A with field a and atom A with field b—all instances of ESD are eliminated, although the amplitudes of $N(t)$ in these subsystems remain relatively low.
As the photon exchange coupling is increased further, with $\kappa>\lambda$ (e.g., $\kappa=5$), ESD in the dynamics of $C(t)$ is removed, since entanglement appears to be nonvanishing (though it shows an irregular behaviour). These observations suggest that the photon exchange interaction actively generates atom-atom entanglement. Simultaneously, the amplitude of field-field entanglement $N(t)$ decreases, which can be attributed to the redistribution of entanglement from the field-field subsystem to the atom-atom subsystem. The frequency of oscillations in the field-field entanglement also increases significantly. Meanwhile, entanglement between atom A and field a diminishes and becomes nearly zero at certain times. The entanglement between atom A and field b does not change considerably, although it vanishes whenever atom-atom and field-field entanglements reach their maxima.
When the photon exchange strength is raised to $\kappa=10$, placing the system in the regime $\kappa\gg\lambda$, the photon exchange interaction dominates over the atom-field coupling. In this scenario, all instances of ESD in the atom-atom entanglement dynamics are eliminated, and $C(t)$ increases significantly in magnitude. This enhancement in atom-atom entanglement corresponds to a marked decrease in field-field entanglement $N(t)$, indicating a transfer of entanglement from the field-field subsystem to the atom-atom subsystem. Furthermore, $N(t)$ for both atom A–field a and atom A–field b subsystems becomes negligible throughout most of the evolution, aside from some small initial peaks.
In summary, strong photon exchange coupling between cavities suppresses ESD in both atom-atom and field-field entanglement dynamics. While the amplitude of $C(t)$ increases notably with increasing $\kappa$, the amplitude of field-field entanglement $N(t)$ diminishes, illustrating a clear transfer of entanglement from the fields to the atoms as photon exchange becomes more dominant.
@@
Figure 7 illustrates the effects of photon exchange interaction on SCTS in the double Jaynes–Cummings model (DJCM). It is observed that the maximum value of $N(t)$ for the field-field subsystem increases by nearly three times due to the photon exchange interaction. For $\kappa=0.1$, the lengths of the entanglement sudden death (ESD) intervals in the atom-atom entanglement increase, and several smaller peaks in the dynamics disappear. The entanglement between atom A and field $a$ reduces significantly, and at certain times, $N(t)$ almost vanishes. In contrast, for the atom A–field $b$ subsystem, the ESDs are removed from the dynamics of $N(t)$. At $\kappa=1$, the ESD duration in $C(t)$ becomes longer as the minor peaks in the dynamics are suppressed. The field-field entanglement $N(t)$ rises sharply and then decreases, although its maximum value becomes lower than that in the $\kappa=0.1$ case. In the atom A–field $a$ subsystem, the increase in $\kappa$ induces ESD, but no significant change is observed in the atom A–field $b$ entanglement.
As $\kappa$ increases further to $5$, the entanglement dynamics of $C(t)$ undergo a drastic transformation. All ESDs are eliminated, and the value of $C(t)$ increases significantly at all times. For the field-field subsystem, the maximum value of $N(t)$ increases, and the entanglement dynamics become highly oscillatory. In both atom A–field $a$ and atom A–field $b$ subsystems, $N(t)$ exhibits initial ESD, but subsequently remains nonzero with very low amplitude. In the regime where $\kappa\gg 1$ (e.g., $\kappa=10$), $C(t)$ increases further and maintains a high value throughout the evolution. Interestingly, the dynamics of $N(t)$ for the field-field entanglement reveal the formation of entanglement patterns, with significantly enhanced amplitudes. Meanwhile, the entanglement between atom A and both fields a and b vanishes completely.
From these observations, it can be concluded that photon exchange interaction primarily generates entanglement in the atom-atom and field-field subsystems. In the case of IDDJCM, we observe that $C(t)$ is on average higher while $N(t)$ for the field-field entanglement is on average lower with respect to DJCM. Conversely, in the standard DJCM, both $C(t)$ and field-field $N(t)$ increase with $\kappa$. In both models, however, the entanglement between atom A and fields a and b vanishes as the photon exchange coupling $\kappa$ increases.
@@ 5 Effects of dipole-dipole interaction between the two atoms on entanglement
Dipole-dipole interaction introduces a direct coupling between the atoms, independent of the field. It adds new energy terms (e.g., exchange of atomic excitation) that interfere with the field-mediated evolution. The resulting energy spectrum becomes anharmonic, leading to incommensurate Rabi frequencies and dephasing.
@@
The effects of dipole–dipole interaction (DDI) between the atoms on the entanglement dynamics for the intensity-dependent double Jaynes–Cummings model (IDDJCM) are presented in Fig. 7. The DDI strength is denoted by $g_{d}$. For lower value such as $g_{d}=0.1$, the entanglement dynamics remain nearly unchanged. This is expected, as weak DDI has minimal influence on the coherent structure of the atom-field subsystems.
However, as the interaction strength increases to $g_{d}=5$, notable changes emerge. The field–field entanglement becomes significantly enhanced in amplitude, though the characteristic oscillations disappear, indicating a shift toward more static field correlations. This reflects the fact that strong DDI enables more effective entanglement transfer between nonlocal subsystems, reducing oscillatory exchange with the local atom–field modes. At $g_{d}=10$, the impact is even more pronounced. All entanglement sudden deaths (ESDs) are removed from the atom–atom concurrence $C(t)$, and while oscillations still occur initially, they gradually damp out over time. The field–field entanglement dynamics stabilize, halting the downward decay typically seen at weaker couplings. In contrast, the atom–field entanglements decay and nearly vanish. This behavior can be interpreted as a consequence of monogamy of entanglement: the strong nonlocal atomic interaction dominates the system dynamics and draws entanglement away from the local atom–field pairs.
The effects of DDI on the standard DJCM with SCTS are shown in Fig. 8. For $g_{d}=0.1$, all subsystems show dynamics similar to the uncoupled case, except the field–field entanglement, which rises in amplitude over time and eliminates ESDs. At $g_{d}=1.0$, the peaks of $C(t)$ are significantly reduced, although the ESD durations remain unchanged. Meanwhile, the field–field entanglement increases further and fully eliminates its own ESDs, suggesting a robust buildup of nonlocal photonic correlations. The amplitude of the atom A–field a entanglement $N(t)$ decreases, whereas $N(t)$ for atom A–field b remains largely unaffected.
For $g_{d}=5$, new peaks emerge in $C(t)$, indicating the onset of additional coherent excitation pathways enabled by the DDI. At $g_{d}=10$, all ESDs in the atom–atom subsystem are fully removed, and the field–field entanglement remains persistently high. However, the atom–field entanglements for both atoms are almost entirely suppressed.
These results highlight the physical role of the dipole–dipole interaction as a mechanism that redistributes entanglement from local (atom–field) to nonlocal (atom–atom and field–field) subsystems. As the DDI strength increases, the system transitions from local coherence-dominated dynamics to a regime where nonlocal correlations dominate, demonstrating a clear manifestation of entanglement monogamy and redistribution.
@@ 6 Effects of spin-spin Ising interaction between the two atoms on entanglement
In the present section, we study the effects of Ising-type interaction between the two atoms on the entanglement dynamics. The Hamiltonian with Ising-type interaction for IDDJCM and DJCM can be written as follows:
@@
The influence of Ising-type interaction on the entanglement dynamics for the intensity-dependent double Jaynes–Cummings model (IDDJCM) is illustrated in Fig. 9. Notably, the behavior observed here differs significantly from the corresponding dynamics in the standard DJCM for squeezed coherent states (SCS) and Glauber–Lachs (G–L) states[59]. As shown in Figs. 9(a) and 9(b), for small values of the Ising interaction strength, namely $J_{z}=0.1$ and $0.3$, the entanglement dynamics of the atom–atom subsystem, $C(t)$, and the atom–field entanglement, $N(t)$, remain nearly indistinguishable from the case with $J_{z}=0$ (see Figs. 1 and 9). However, subtle effects are observed in the cross-coupled subsystems. Specifically, the entanglement between atom A and field $b$ shows a reduction in the amplitude of $N(t)$, along with the persistence of entanglement sudden deaths (ESDs). In contrast, the field aa–field $b$ entanglement shows notable enhancement: the amplitude increases and ESDs are eliminated, suggesting that the Ising interaction facilitates entanglement redistribution into the field–field sector.
As the Ising interaction strength increases to $J_{z}=0.7$, new features emerge in the atom–atom entanglement dynamics. Small secondary peaks begin to appear in $C(t)$, accompanied by a reduction in the overall amplitude. For the atom A–field $a$ subsystem, $N(t)$ shows a mild decrease in amplitude, while ESDs are gradually removed. Interestingly, the atom A–field $b$ dynamics remain largely unaffected. However, the field–field entanglement continues to grow in amplitude and shows no signs of ESD, indicating stronger nonlocal correlations mediated by the Ising interaction. At $J_{z}=1.0$, the small peaks in $C(t)$ appear earlier in the evolution and grow over time, whereas the initial large peak diminishes, resulting in a net reduction in the duration of ESDs. In atom–field entanglements, the peak heights decrease for atom A–field $a$, but increase for atom A–field $b$ and field aa–field $b$ subsystems.
@@
The effects of Ising interaction on entanglements in DCJM are depicted in Fig. 10. In this case also, we observe that for small values of $J_{z}$ (0.1 and 0.3), $C(t)$ and $N(t)$ for all subsystems do not change considerably. Only field-field entanglement increases slightly for $J_{z}=0.3$. For $J_{z}=0.7$, length of ESD for $C(t)$ increases slightly, as the small bump disappears from the dynamics. The amplitudes of atom-field entanglement do not change noticeably; however, the length of ESD decreases for atom A-field b entanglement. For $J_{z}=1.0$, two small peaks appear in the dynamics of $C(t)$ which makes the length of ESDs smaller. All other entanglement dynamics remain almost unchanged.
These observations suggest that the Ising-type interaction introduces subtle redistributions of entanglement among the subsystems, particularly enhancing nonlocal correlations in the field sector. However, it is not significantly effective in eliminating ESDs in the atom–atom or cross atom–field subsystems. The nonlinearity introduced by the Ising term, while capable of modifying local coherence and interference effects, does not dominate the entanglement evolution unless its strength is significantly increased.